OpenStudy (anonymous):

Solve the equation: (3x - 2)(7x + 1) = 0.

OpenStudy (callisto):

(3x - 2)(7x + 1) = 0 Put (3x - 2) =0 or (7x + 1) = 0 x = 2/3 x=-1/7 actually do you understand how to deal with such kind of questions?

OpenStudy (anonymous):

I don;t :(

OpenStudy (callisto):

Give you an easy one to try x (5) = 0, what is x?

OpenStudy (anonymous):

0

OpenStudy (callisto):

How can you get the answer?

OpenStudy (anonymous):

Multiplying something by zero always results in a zero.

OpenStudy (callisto):

Yup.. that's the same when you consider the xy=0 What would you think if xy=0? what is the possible value for x or y?

OpenStudy (callisto):

Hint: remember what you've said just now 'Multiplying something by zero always results in a zero.'

OpenStudy (anonymous):

I'm not exactly sure how to work the double letters unfortunately. I know if you multiple one by zero it gets zero? I think.

OpenStudy (callisto):

Yup.. you're quite close to the point.. For the product of 2 unknowns, the either one should be =0 to give 0 as the result (of course both of them can be 0), Do you understand so far?

OpenStudy (anonymous):

Yeah, I understand so far. Until numbers get added into the situation which makes it twice as difficult.

OpenStudy (callisto):

When you see something like (3x - 2)(7x + 1) = 0. Just consider (3x - 2) =y and (7x + 1)=z So yz=0 Understand so far?

OpenStudy (anonymous):

Sort of.

OpenStudy (callisto):

So, either y = 0 or z=0 can make the product =0, understand?

OpenStudy (anonymous):

Now it makes more sense.

OpenStudy (callisto):

So remember you let y=(3x - 2) and z=(7x + 1) Put it back to y = 0 or z=0 respectively, and solve for it separately, understand?

OpenStudy (anonymous):

That part kind of confuses me the final solve.

OpenStudy (callisto):

Hmmm... which part you don't understand?

OpenStudy (anonymous):

So remember you let y=(3x - 2) and z=(7x + 1) Put it back to y = 0 or z=0 respectively, and solve for it separately, understand?

OpenStudy (callisto):

So remember you let y=(3x - 2) and z=(7x + 1) <- understand?

OpenStudy (anonymous):

Yes I understand that part.

OpenStudy (callisto):

either y = 0 or z=0 can make the product =0, understand <- understand?

OpenStudy (anonymous):

That's where I get confused.

OpenStudy (callisto):

yz=0 either y = 0 or z=0 can make the product =0, understand?

OpenStudy (anonymous):

Yeah

OpenStudy (callisto):

so y= 3x-2 =0 or z=7x+1=0, understand?

OpenStudy (anonymous):



OpenStudy (callisto):

So 3x-2 =0 or 7x+1=0 and solve the 2 equations, understand?

OpenStudy (anonymous):

I don't know how to solve them both together.

OpenStudy (callisto):

3x-2 =0 x = 2/3 or 7x+1 =0 x=-1/7 Just solve it separately